OpenStudy (anonymous):

Find two consecutive even integers such that the square of the smaller is 10 more than the larger.

OpenStudy (cwrw238):

numbers are n and n + 2 n^2 = n + 12 n^2 - n - 12 = 0 (n - 4)(n+3) = 0 n = 4 2 numbers are 4 and 6

OpenStudy (cwrw238):

do you follow that alright?